# Transcorrelated Theory with Pseudopotentials

### v1

@@ 3 Calculations
The transcorrelated, second-quantized Hamiltonians are calculated with an in-house code TCHINT 28, that is based on the version used in 7. TCHINT calculates transcorrelated second-quantized Hamiltonians with a Jastrow factor and the molecular orbitals as inputs. It is parallelized with MPI, uses the BLAS and LAPACK libraries for matrix operations, and is written in Fortran. The inclusion of the features necessary for the PP commutator evaluation within TCHINT has been an important part of this work. The PP commutators evaluations are calculated as integrals over real space, and for each real space grid point within the influence of a pseudopotential we define a spherical grid of points, centered around the atom, to calculate spherical projections. This introduces additional Jastrow evaluations and a possible load imbalance problem for MPI parallelization. To mitigate this, we use vectorized Jastrow evaluations and a load balancing scheme that distributes the grid points evenly among the MPI processes.
@@
The treatment of the $3$-body terms when constructing the transcorrelated Hamiltonian is done under the xTC approximation 14. In this approximation, the last term of Eq. 10 containing $L$ is reorganized within the generalized normal ordering scheme. This leads to modifications in the $1$-,$2$-, and $3$-body terms of the transcorrelated Hamiltonian. The contributions of the $3$-body terms under the generalized normal ordering to the $1$- and $2$-body terms are evaluated by contracting the $3$-body terms with the reduced $1$-body density matrix of the Hartree-Fock wave function (with the exception of using FCI density matrix in the N2 dissociation curve calculations, see Sec. 4.4.1). The remaining 3-body terms are neglected in xTC.
@@
The atomization energies are calculated both with and without transcorrelation with the coupled cluster using singles, doubles and perturbative and full triples (CCSD(T) and CCSDT) using the ElemCo.jl package 29. If transcorrelation is used, we add a prefix xTC- to the method name. The similarity transformation of the Hamiltonian using the Jastrow factor leads to a non-Hermitian Hamiltonian with a non-diagonal Fock matrix. Consequently, standard non-iterative perturbative methods for CCSD(T) are not directly applicable to the transcorrelated Hamiltonian. Thus for transcorrelated CCSD(T) we do the calculations with a pseudocanonical $\Lambda$CCSD(T) approach using bi-orthogonal orbitals 30. To simplify notation we will call it xTC-CCSD(T).
@@
The atomization energies are also calculated with CCSD(T)-F12 using the Molpro package 31, 32, 33 for comparison. Both all-electron and PP F12 results are calculated to assess the effect of the PPs on the results. In all-electron calculations we use the standard AVXZ family of basis sets with AVXZ-MP2Fit and VXZ-JKFit auxiliary basis sets 34. In PP calculations we use the augmented basis sets fitted for PPs 21, 22. As the auxiliary basis sets with PPs we use mp2-fitted QZVPP/MP2Fit and TZVPP/MP2Fit basis sets.
@@
For the dissociation curves, we use the full configuration interaction quantum Monte Carlo method (FCIQMC) 35 with the NECI package 36 both with and without transcorrelated Hamiltonians. The initiator approximation 35 with an initiator treshold of $3$ is used in the FCIQMC calculations. The walker number for each calculation was increased by a factor of $5$ until the energy was converged to within $1$ mHa. We compare the results with MRCI-f12 calculations, done with the Molpro package 31, 32, 33. The Davidson correction 37 is used in the MRCI-F12 calculations.
@@
When presenting the total energies of molecules with PPs, we show the results relative to an energy $E_{CBS}^{TZ-QZ}$, evaluated as the sum of Hartree-Fock energy in the PVQZ (ccECPs) or AVQZ (eCEPPs) basis set, $E_{HF}^{QZ}$, and the estimate of the complete basis set limit (CBS) of CCSD(T) correlation energy. This estimate is obtained from the PVTZ and PVQZ (ECP) or AVTZ and AVQZ (eCEPP) correlation energies $E^{Corr}_{TZ}(\mathrm{CCSD(T)})$ and $E^{Corr}_{QZ}(\mathrm{CCSD(T)})$ with a linear extrapolation, so that
$E_{CBS}^{TZ-QZ}(\mathrm{CCSD(T)})=E_{HF}^{QZ}+\frac{3^{3}E^{Corr}_{TZ}(\mathrm{CCSD(T)})-4^{3}E^{Corr}_{QZ}((\mathrm{CCSD(T)}))}{3^{3}-4^{3}}.$ (13)
@@ 4.1 Variances in VMC optimization
Table 1 shows the variance of the reference energy in Hartrees for the first row elements Be-F, and for a set of first row molecules, obtained from eCEPP, ccECP, and all-electron (AE) calculations. For the results, we sampled the reference Hartree-Fock wave function with the Metropolis algorithm and evaluated an estimate of the variance of the obtained configurations together with the optimized Jastrow factor. The percentages after the PP variances show the ratio of PP values against the all-electron variances. The variance is significantly reduced when using PPs as compared to AE calculations. The variance reduction is greater for the heavier atoms.
@@
Table 2: Non-local cutoff radius values for various atoms, in atomic units, for eCEPPs and ccECPs. The cutoff is defined as the radius above which the PP radial functions are less than $10^{-6}$ Ha.
@@ 4.2.1 Analysis of the degree of PP commutators
In Fig. 2 we show the xTC-CCSD(T)(PP-$n$) total energies of the first row elements Be-F, along with their ionised states, as a function of $n$, the degree of PP commutators evaluated. Results with both eCEPPs (a) and ccECPs (b) are shown. The results are calculated with the quadruple-zeta basis set. The zero in the y-axis refers to $E_{CBS}^{TZ-QZ}(\mathrm{CCSD(T)})$ (see Eq. (13)).
@@
Figure 3 shows the ionisation energies of the first row elements Be-F, obtained with CCSD(T), xTC-CCSD(T)(PP-00), xTC-CCSD(T)(PP-$1$), and xTC-CCSD(T)(PP-$2$) using both eCEPPs (a) and ccECPs (b). The results are shown as a function of the basis set. The ionisation energies are shown against experimental values 38.
CCSD(T) ionisation energies reach chemical accuracy for Be, C, and N with ccECPs and PVQZ basis set. With eCEPPs and AVQZ basis only Be ionisation energy with CCSD(T) is chemically accurate, but B, C, and N are close to chemical accuracy.
@@
(a) eCEPPs:
(b) ccECPs:
Figure 3: Ionisation energies (IPs) $E_{i}=E_{atom}-E_{ion}$ for the first row elements, using (a) eCEPPs and (b) ccECPs. The energies are presented as the discrepancy with the experimental ionisation energies 38, so that the presented energies are $E_{method}^{basis}-E_{exp}$. The X-axis shows the xTC-CCSD(T)(PP-$n$) ionisation energies with $n=0,1,2$ (in 2nd, 3rd, and 4th column, respectively). The 1st column shows CCSD(T) ionisation energies. The basis sets used are the AVXZ series for eCEPPs and the PVXZ series for ccECPs, where X is either D, T, or Q in the 1st, 2nd, and 3rd row, respectively. The grey shaded region denotes chemical accuracy. The red shading denotes a sum of gaussians, centered at each data point, with a width set such that equidistant gaussians in the presented scale would overlap at $95$ confidence.
@@ 4.3.1 Total transcorrelated energies
Figure 4: Energies of the molecules CN, CO, CF, N2, O2, F2, H2O, and CO2 with eCEPPs (a) and ccECPs (b) as a function of the basis set. The energies are relative to CCSD(T) CBS energies as $\Delta E=E-E_{CBS}^{TZ-QZ}(\mathrm{CCSD(T)})$, see Eq. (13). CCSD(T) (green), xTC-CCSD(T)(PP-00) (yellow), and xTC-CCSD(T)(PP-$2$) (violet) energies are shown.
Figure 4 shows the energies of the molecules CN, CO, CF, N2, O2, F2, H2O, and CO2, obtained with eCEPPs (a) and ccECPs (b) with CCSD(T) and xTC-CCSD(T)(PP-$n$) methods with $n=0$ and $n=2$. The energies are displayed relative to $E_{CBS}^{TZ-QZ}(\mathrm{CCSD(T)})$. The results are shown as a function of the basis set.
@@ 4.3.2 F12 atomization energies
Figure 5 shows the atomization energies of the molecules CN, CO, CF, N2, O2, F2, H2O, and CO2, calculated with CCSD(T)-F12. The results are shown relative to the HEAT database 24. The results are obtained with the AVXZ basis sets, with X=D,T,Q. The results are shown for all-electron, eCEPP, and ccECP calculations.
@@
Figure 5: Atomization energies of the molecules CN, CO, CF, N2, O2, F2, H2O, and CO2, evaluated as $E_{at}=\sum E_{atom}-E_{mol}$, calculated with CCSD(T)-F12 and MRCI-F12 methods. The energies are presented relative to the results in the HEAT database, so that the presented energies are $E_{HEAT}-E_{at}$. Results are obtained using AVXZ basis sets, with X=D,T,Q (see labels on right). Results from all-electron (top), eCEPP (middle), and ccECP (bottom) calculations are shown. The grey shaded region denotes chemical accuracy. The red shading represents a sum of gaussians centered at the atomization energy discrepancies of each molecule, with sigmas set so that equidistantly placed gaussians over the interval between maximum and minimum of any calculation with a given core treatment (AE, eCEPP, ccECP) would have nearest-neighbour distance of $4$ sigma.
@@ 4.3.3 Atomization energies with transcorrelation
We calculated the atomization energies of the molecules CN, CO, CF, N2, O2, F2, H2O, and CO2, using eCEPPs and ccECPs with AVXZ basis sets (Figures 6 and 7). The results are presented as discrepancies relative to the HEAT values 24, employing CCSD(T) and CCSDT 29. Both methods were applied in three variants: non-TC, xTC without PP commutator evaluations, and xTC with two PP commutator evaluations.
@@
Figure 6: Atomization energies of a test set of molecules, evaluated with eCEPPs as $E_{at}=\sum E_{atom}-E_{mol}$. The energies are presented relative to the results in the HEAT database, so that the presented energies are $E_{HEAT}-E_{at}$. Calculations are shown without TC (left column),and with xTC both without and with 2 PP commutator evaluations. The grey shaded region denotes chemical accuracy. The red shading represents a sum of gaussians centered at the atomization energy discrepancies of each molecule, with sigmas set so that equidistantly placed gaussians over the interval between maximum and minimum of a given level of theory would have nearest-neighbour distance of $4$ sigma.
@@
Table 5 summarizes the mean absolute errors (MAE) and root mean square errors (RMS) for all combinations of PP type, method, basis set, and theory. In terms of MAE and RMS, Non-TC methods remain far from chemical accuracy even with the AVQZ basis set. In contrast, transcorrelated methods with eCEPPs, AVQZ basis sets, and two PP commutator evaluations achieve chemical accuracy across all studied methods. With ccECPs, xTC methods with two PP commutator evaluations are within or very close to chemical accuracy with AVQZ basis, and within chemical accuracy using the AVTZ basis set. The fact that we see best results with AVTZ basis, and not with AVQZ basis, when using ccECPs both with F12 and transcorrelated methods hints that the non-monotonic increase in accuracy with basis set resolution is a feature of the ccECPs.

### v2

@@ 3.1 Evaluation of transcorrelated Hamiltonian with pseudopotentials
The transcorrelated second-quantized Hamiltonians are calculated with an in-house code TCHINT 26, that is based on the version used in 7. TCHINT calculates transcorrelated second-quantized Hamiltonians with a Jastrow factor and the molecular orbitals as inputs. It is parallelized with MPI, uses the BLAS and LAPACK libraries for matrix operations, and is written in Fortran. The inclusion of the features necessary for the PP commutator evaluation within TCHINT has been an important part of this work.
The elements $P_{rs}^{pq}$ in Eq. (11) are obtained with numerical integration over real-space grid points. The operation of $\hat{H}^{\text{PP}}_{\text{en}}$ in terms $\Pi$ and $\Gamma$ is evaluated according to Eq. (5), by discretizing the spherical integration, so that
$\hat{V}_{\text{eff}}(r_{i})f(\mathbf{r}_{i})=\sum_{l=0}^{l_{\text{max}}}V_{l}(r_{i})Y_{l0}(\Omega_{\mathbf{r}_{i}})\sum_{i=1}^{N_{s}}f(\mathbf{r}_{i}{{}^{\prime}})Y_{l0}(\Omega_{\mathbf{r}_{i}{{}^{\prime}}}),$ (13)
where $\left|\mathbf{r}_{i}{{}^{\prime}}\right|=\left|\mathbf{r}_{i}\right|=r_{i}$ and f is either $\phi_{r}(\mathbf{r}_{1})$, $J_{2}(\mathbf{r}_{1},\mathbf{r}_{2})\phi_{r}(\mathbf{r}_{1})$, or $J_{2}(\mathbf{r}_{1},\mathbf{r}_{2})^{2}\phi_{r}(\mathbf{r}_{1})$. The spherical grid points are chosen as the vertices of an icosahedron. The points are obtained by setting them on unit sphere scaled with $r_{i}$, so that [±$[\pm a,\pm b,0]$, [±$[\pm b,\pm a,0]$, and $[0,\pm a,\pm b]$ are the grid points $\mathbf{r}_{i}{{}^{\prime}}$ on the unit sphere, with $a=r_{i}/\sqrt{1+\phi^{2}}$ and $b=\phi r_{i}/\sqrt{1+\phi^{2}}$, where $\phi=(1+\sqrt{5})/2$ is the golden ratio. Hence $N_{s}=12$. This is a Lebedev grid capable of exact spherical integration of functions that have up to $l=5$ components. Tests on denser spherical grids did not change the results significantly. To mitigate the bias by the orientation of the spherical grid we applied random rotations to the spherical grid points in each spherical projection.
For a number of grid points $N_{\text{ECP}}$ under pseudopotential influence, the numbers of additional Jastrow factor and orbital evaluations due to presence of pseudopotentials are $N_{\text{g}}N_{\text{ECP}}N_{\text{s}}$ and $N_{\text{ECP}}N_{\text{orb}}N_{\text{s}}$, respectively, where $N_{\text{g}}$ is the total number of grid points and $N_{\text{orb}}$ is the number of orbitals. The additional jastrow evaluations thus add a $N_{\text{ECP}}N_{s}/N_{g}$ prefactor to the computational scaling of the Jastrow evaluations 7, which is computationally most expensive part of the calculation. To mitigate the additional computational cost and load imbalance, we use vectorized Jastrow evaluations and a load balancing scheme that distributes the grid points evenly among the MPI processes. Orbitals at the spherical grid points are evaluated at the beginning of the calculation, and the memory requirements are increased because of ECPs by additional $N_{\text{ECP}}N_{\text{orb}}N_{\text{s}}$ orbital floating point numbers as opposed to the $N_{\text{g}}N_{\text{orb}}$ floating point numbers required for all-electron calculations.
Table 1 shows a list of the number of grid points under pseudopotential influence for a number of systems studied in this work. For each system, the number of grid points is roughly 55% of the total number of grid points.
Table 1: Total and ECP-influenced grid points for each system and pseudopotential.
System Total eCEPP ccECP
C 18120 10570 10268
N 18120 10570 9966
O 18120 10268 9362
F 18120 9966 9060
N2 36196 22792 20810
O2 36198 21380 19058
F2 36202 20316 18324
CN 36196 22445 21045
CO 36196 22022 20313
CF 36198 21299 19767
H2O 33484 20066 9612
CO2 54056 33384 30250
@@
The treatment of the $3$-body terms when constructing the transcorrelated Hamiltonian is done under the xTC approximation 15. In this approximation, the last term of Eq. 10 containing $L$ is reorganized within the generalized normal ordering scheme. This leads to modifications in the $1$-,$2$-, and $3$-body terms of the transcorrelated Hamiltonian. The contributions of the $3$-body terms under the generalized normal ordering to the $1$- and $2$-body terms are evaluated by contracting the $3$-body terms with the reduced $1$-body density matrix of the Hartree-Fock wave function (with the exception of using FCI density matrix in the N2 dissociation curve calculations. The remaining 3-body terms are neglected in xTC.
3.2 General computational details
In this work we study the use of pseudopotentials with the transcorrelated Hamiltonian in atoms Be, B, C, N, O, and F, as well as their $+1$ ions. We also study the total and atomization energies of molecules CN, CO, CF, N2, O2, F2, H2O and CO2. We use the aug-cc-pVD/T/QZ basis sets (AVXZ, with X=D,T,Q), optimized individually for each pseudopotential22, 23. Ionization energies with ccECPs are evaluated with non-augmented cc-pVD/T/QZ basis sets (PVXZ, with X=D,T,Q). .
The geometries of the molecules are taken from those in the HEAT database 27. The Hartree-Fock (HF) calculations are done with the PYSCF code 28. The Jastrow factors used are of Drummond-Towler-Needs type 25 and are optimised with respect to the variance of the Hartree-Fock wave function using the VMC method of the CASINO package 29. The Jastrow factors are optimized separately for each system, basis set, and pseudopotential combination. The cutoffs used for the $u$, $\chi$, and $f$ terms were $4.5$, $4$, and $4$, respectively. We provide the optimized Jastrow factors in the Supplementary Material.
@@ 3.2 General computational details
The atomization energies are calculated both with and without transcorrelation with the coupled cluster using singles, doubles and perturbative and full triples (CCSD(T) and CCSDT) using the ElemCo.jl package 30. If transcorrelation is used, we add a prefix xTC- to the method name. The similarity transformation of the Hamiltonian using the Jastrow factor leads to a non-Hermitian Hamiltonian with a non-diagonal Fock matrix. Consequently, standard non-iterative perturbative methods for CCSD(T) are not directly applicable to the transcorrelated Hamiltonian. Thus for transcorrelated CCSD(T) we do the calculations with a pseudocanonical $\Lambda$CCSD(T) approach using bi-orthogonal orbitals 31. To simplify notation we will call it xTC-CCSD(T).
@@
The atomization energies are also calculated with CCSD(T)-F12 using the Molpro package 32, 33, 34 for comparison. Both all-electron and PP F12 results are calculated to assess the effect of the PPs on the results. In all-electron calculations we use the standard AVXZ family of basis sets with AVXZ-MP2Fit and VXZ-JKFit auxiliary basis sets 35. In PP calculations we use the augmented basis sets fitted for PPs 22, 23. As the auxiliary basis sets with PPs we use mp2-fitted QZVPP/MP2Fit and TZVPP/MP2Fit basis sets.
@@
For the dissociation curves, we use the full configuration interaction quantum Monte Carlo method (FCIQMC) 36 with the NECI package 37 both with and without transcorrelated Hamiltonians. The initiator approximation 36 with an initiator threshold of $3$ is used in the FCIQMC calculations. The walker number for each calculation was increased by a factor of $5$ until the energy was converged to within $1$ mHa. We compare the results with MRCI-F12 calculations, done with the Molpro package 32, 33, 34. The Davidson correction 38 is used in the MRCI-F12 calculations.
3.3 Notation
@@
[…]
$E_{\text{CBS}}^{\text{TZ-QZ}}(\text{CCSD(T)})=E_{\text{HF}}^{\text{QZ}}+\frac{3^{3}E^{\text{Corr}}_{\text{TZ}}(\text{CCSD(T)})-4^{3}E^{\text{Corr}}_{\text{QZ}}((\text{CCSD(T)}))}{3^{3}-4^{3}}.$ (14)
@@ 4.1 Variances in VMC optimization
Table 2 shows the variance of the reference energy in Hartrees for the first row elements Be-F, and for a set of first row molecules, obtained from eCEPP, ccECP, and all-electron (AE) calculations. For the results, we sampled the reference Hartree-Fock wave function with the Metropolis algorithm and evaluated an estimate of the variance of the obtained configurations together with the optimized Jastrow factor. The percentages after the PP variances show the ratio of PP values against the all-electron variances. The variance is significantly reduced when using PPs as compared to AE calculations. The variance reduction is greater for the heavier atoms. For the atoms, the variance reduction seems to obey roughly $1/N_{v}$ dependence, where $N_{v}$ is the number of valence electrons.
@@
Table 3: Non-local cutoff radius values (Bohr) for various atoms, in atomic units, for eCEPPs and ccECPs. The cutoff is defined as the radius above which the PP radial functions are less than $10^{-6}$ Ha.
@@
4.2 Analysis of the transcorrelated integrals
For the systems studied in this work, we have investigated statistical parameters of the off-diagonal values of $V$, $K$, and $P$ tensors, as well as the full TC Hamiltonian. The minima, maximas, mean values, and Frobenius norms of the tensors are shown in Fig. 2 for both ccECPs and eCEPPs. The values are shown in Hartrees. Frobenius norm is defined as $F=\sqrt{\sum_{ij}A_{ij}^{2}}$ for a matrix $A$.
The data shows that the largest values and overall weight of the off-diagonals are in the $V$ tensor, i.e., the non-transcorrelated Hamiltonian, with the full TC Hamiltonian having slightly smaller values and overall weight (the Frobenius norm) than $V$. The $K$ tensor of the kinetic energy operator commutators has much smaller weight compared to the full Hamiltonian, and the statistical parameters of the $P$ tensor show that the pseudopotential commutators introduce only a slight correction to the TC Hamiltonian.
This data shows that pseudopotential commutators introduce generally small but non-negligible contributions.
Figure 2: Parameters of the off-diagonal values of $V$ (green), $K$ (blue), $P$ (orange), and $V-K+P$ (red) tensors of Eq. 10 for different systems. The minima (upper left), maximas (upper right), mean values (lower left), and Frobenius norms (lower right) of the tensors are shown in Hartrees. For each stochastic parameter, we show values for each system for both ccECPs and eCEPPs.
@@ 4.3.1 Analysis of the degree of PP commutators
[…]
@@
Figure 4 shows the ionisation energies of the first row elements Be-F, obtained with CCSD(T), xTC-CCSD(T)(PP-00), xTC-CCSD(T)(PP-$1$), and xTC-CCSD(T)(PP-$2$) using both eCEPPs (a) and ccECPs (b). The results are shown as a function of the basis set. The ionisation energies are shown against experimental values 39.
CCSD(T) ionisation energies reach chemical accuracy ($1.6$mHa, $0.04$eV) for Be, C, and N with ccECPs and PVQZ basis set. With eCEPPs and AVQZ basis only Be ionisation energy with CCSD(T) is chemically accurate, but B, C, and N are close to chemical accuracy.
@@
(a) 1st ionisation energies, eCEPPs:
(b) 1st ionisation energies, , ccECPs:
Figure 4: Ionisation energies (IPs) $E_{\text{i}}=E_{\text{atom}}-E_{\text{ion}}$ for the first row elements, using (a) eCEPPs and (b) ccECPs. The energies are presented as the discrepancy with the experimental ionisation energies 39, so that the presented energies are $\Delta E=E_{\text{method}}^{\text{basis}}-E_{\text{exp}}$. The X-axis shows the xTC-CCSD(T)(PP-$n$) ionisation energies with $n=0,1,2$ (in 2nd, 3rd, and 4th column, respectively). The 1st column shows CCSD(T) ionisation energies. The basis sets used are the AVXZ series for eCEPPs and the PVXZ series for ccECPs, where X is either D, T, or Q in the 1st, 2nd, and 3rd row, respectively. The grey shaded region denotes chemical accuracy. The red shading denotes a sum of gaussians, centered at each data point, with a width set such that equidistant gaussians in the presented scale would overlap at $95$ confidence.
@@ 4.4.1 Total transcorrelated energies
[…]
@@ 4.4.2 F12 atomization energies
Figure 6 shows the atomization energies of the molecules CN, CO, CF, N2, O2, F2, H2O, and CO2, calculated with CCSD(T)-F12. The results are shown relative to the HEAT database 27. The results are obtained with the AVXZ basis sets, with X=D,T,Q. The results are shown for all-electron, eCEPP, and ccECP calculations.
@@
Figure 6: Atomization energies of the molecules CN, CO, CF, N2, O2, F2, H2O, and CO2, evaluated as $E_{\text{at}}=\sum E_{\text{atom}}-E_{\text{mol}}$, calculated with CCSD(T)-F12 method. The energies are presented relative to the results in the HEAT database, so that the presented energies are $\Delta E=E_{\text{HEAT}}-E_{\text{at}}$. Results are obtained using AVXZ basis sets, with X=D,T,Q (see labels on right). Results from all-electron (top), eCEPP (middle), and ccECP (bottom) calculations are shown. The grey shaded region denotes chemical accuracy. The red shading represents a sum of gaussians centered at the atomization energy discrepancies of each molecule, with sigmas set so that equidistantly placed gaussians over the interval between maximum and minimum of any calculation with a given core treatment (AE, eCEPP, ccECP) would have nearest-neighbour distance of $4$ sigma.
@@ 4.4.3 Atomization energies with transcorrelation
We calculated the atomization energies of the molecules CN, CO, CF, N2, O2, F2, H2O, and CO2, using eCEPPs and ccECPs with AVXZ basis sets (Figures 7 and 8). The results are presented as discrepancies relative to the HEAT values 27, employing CCSD(T) and CCSDT 30. Both methods were applied in three variants: non-TC, xTC without PP commutator evaluations, and xTC with two PP commutator evaluations.
@@
Figure 7: Atomization energies of a test set of molecules, evaluated with eCEPPs as $E_{\text{at}}=\sum E_{\text{atom}}-E_{\text{mol}}$. The energies are presented relative to the results in the HEAT database, so that the presented energies are $\Delta E=E_{\text{HEAT}}-E_{\text{at}}$. Calculations are shown without TC (left column),and with xTC both without and with 2 PP commutator evaluations ({ method}(0) and { method}(2), respectively, method being either CCSD(T) or CCSDT). The grey shaded region denotes chemical accuracy. The red shading represents a sum of gaussians centered at the atomization energy discrepancies of each molecule, with sigmas set so that equidistantly placed gaussians over the interval between maximum and minimum of a given level of theory would have nearest-neighbour distance of $4$ sigma. To improve presentation and to better compare with CCSD(T)-F12 results, we are only showing the region $-0.1<\Delta E<0.5$eV, which leaves some of the energies of non-transcorrelated methods outside of the range in AVDZ and AVTZ basis sets.
@@
Table 6 summarizes the mean absolute errors (MAE) and root mean square errors (RMS) for all combinations of PP type, method, basis set, and theory. In terms of MAE and RMS, non-transcorrelated methods remain far from chemical accuracy even with the AVQZ basis set. In contrast, transcorrelated methods with eCEPPs, AVQZ basis sets, and two PP commutator evaluations achieve chemical accuracy across all studied methods. With ccECPs, xTC methods with two PP commutator evaluations are within or very close to chemical accuracy with AVQZ basis, and within chemical accuracy using the AVTZ basis set. The fact that we see best results with AVTZ basis, and not with AVQZ basis, when using ccECPs both with F12 and transcorrelated methods hints that the non-monotonic increase in accuracy with basis set resolution is a feature of the ccECPs.